Two dice are rolled. Determine the probability of the following.
Rolling an even number or a number greater than 9

Answer by Alex.33(110)

I'll now assume the "number" in your question to be the sum of the number given by the two dice. (Otherwise it is not possible for a typical dice to roll a number >9)
The total number of possible outcomes is 6*6=36 cases.
a. To get an even sum, the two numbers are supposed to be even+even(3*3=9) or odd+odd(3*3=9)/*this is simple number theory, simply provable*/, with a total of 18 cases under this condition.
P(an even number)=18/36=1/2. /*since this is a classical probability question, I use the classical definition here, although it is actually somewhat non-rigorous in logic. If curious you can visit wikipedia: probability.*/
b. To get a number greaer than 9, there are 3 kinds of cases:
sum=10: 4+6, 5+5, 6+4.
sum=11: 5+6, 6+5.
sum=12: 6+6.
A total of 6 cases.
P(a number greater than 9)=6/36=1/6.